OpenStudy (anonymous):

A local company manufactures netbook computers. Their profit function is given by this equation: y=-x^2+120x+3000. where x is the number of netbooks produced in one day and y is profit (in dollars) the company makes that day. What is the maximum profit that the company can make? (Find the vertex of the parabola)

OpenStudy (danjs):

do you take calc?

OpenStudy (anonymous):

no

OpenStudy (danjs):

dy/dx = -2x + 120 -2x + 120 = 0 x = 60

OpenStudy (danjs):

x= 60, y = 6600 profit max

OpenStudy (anonymous):

So is that the answer?

OpenStudy (danjs):

yeah, the vertex is at (60,6600) and the parabola faces down

OpenStudy (anonymous):

ok thanks!! And can i ask you one more question?

OpenStudy (jhannybean):

Or no.

OpenStudy (danjs):

im out.. cyas

OpenStudy (jhannybean):

Sorry, your equation is \(y=-x^2+120x+3000\) and your vertex would be at \(x-\dfrac{-b}{2a}\)

OpenStudy (jhannybean):

\[x=\frac{-b}{2a}\]

OpenStudy (jhannybean):

If you have not learned calculus, then you simply need to find the x coordinate of your vertex, use that to find your y-coordinate of the vertex, and you will have the coordinate points for your vertex.

OpenStudy (anonymous):

so it would be X=-(-1) over 2(-1)?

OpenStudy (jhannybean):

Let's check.\[x=-\frac{b}{2a} = \frac{-120}{-2(1)} =~?\]

OpenStudy (anonymous):

ohh forgot the -120 bit

OpenStudy (jhannybean):

Analyze your formula :\[y=\color{red}{-x^2}\color{blue}{+120x}\color{green}{+3000}\]\[y=\color{red}{a}x^2\color{blue}{+b}x\color{green}{+c}\]

OpenStudy (anonymous):

ok so we a which is -1 and b=120, c=3000

OpenStudy (jhannybean):

Here im showing you that your equation matches the general formula for a quadratic equation.

OpenStudy (jhannybean):

yes! :)

ganeshie8 (ganeshie8):

completing the square would be more easy here

OpenStudy (anonymous):

yes!! ok

OpenStudy (jhannybean):

Let's try a different method this time @ganeshie8 hahaah

ganeshie8 (ganeshie8):

you had breakup wid completing the square is it hmm

OpenStudy (jhannybean):

Yeah I agree it would be easier, but first a person needs to learn which values match up to the values in a quadratic :P

OpenStudy (jhannybean):

Yeah I diid lol

OpenStudy (jhannybean):

So what would be your x-value, @Lovelucy19 ?

ganeshie8 (ganeshie8):

why do you need to learn that for completing the square ?

OpenStudy (anonymous):

60?

OpenStudy (jhannybean):

You don't, it's just a different method for finding the vertex I think.I'll show her compelting the square afterwards :P

ganeshie8 (ganeshie8):

OK - vertex form gives geometric intuition of the profit function, this -b/2a thingy is more like memorizing and algebra

OpenStudy (jhannybean):

Good, \(x=60\). Now we can write our equation, \(y=-x^2+120x+3000\) as \(f(x)=-x^2+120x+3000\) therefore,since we found our \(x\) value, we can plug that in for our function, \(f(60) = -(60)^2 +120(6)+3000\)

OpenStudy (anonymous):

Question why don't we put anything on the C part?

OpenStudy (jhannybean):

Is there an x in the c part of our function?

OpenStudy (jhannybean):

if there is not, then we cannot put anything there :P

OpenStudy (anonymous):

ohh ok i was getting a little confused, but im good now

OpenStudy (jhannybean):

Oh, good. :) So what do we get for f(60)?

OpenStudy (anonymous):

isn't that just 60 still or would that be -60?

OpenStudy (jhannybean):

For the first term? \(-(60)^2 = -(60 \cdot 60) =~?\)

OpenStudy (anonymous):

3,600

OpenStudy (jhannybean):

And then you multiply the negative to it.

OpenStudy (anonymous):

-3600

OpenStudy (jhannybean):

remember, if the negative is outside, \(-(60)^2\) vs \((-60)^2\), the first one will give you \(-3600\) and the second one will give you 3600. that is because you multiply a -1 to a function that you first square, the first one, and for the second one, you are doing this: \((-60) \cdot (-60)\)

OpenStudy (jhannybean):

If that makes any sense, so you have -3600 for the first term . then add it to \(120 \cdot 6\). What does that give you?

OpenStudy (anonymous):

720

OpenStudy (jhannybean):

What I meant is: -3600 + (120 * 6) =?

OpenStudy (jhannybean):

and yes, 120 * 6 = 720 \(\checkmark\)

OpenStudy (anonymous):

ok so itwouuld be -2880

OpenStudy (jhannybean):

Yes :)

OpenStudy (jhannybean):

-2880 +3000?

OpenStudy (anonymous):

120

OpenStudy (jhannybean):

Yes ) so y=120

OpenStudy (jhannybean):

Therefore our vertex is \(\boxed{(60,120)}\)

OpenStudy (anonymous):

Thanks so much for your help with this I honestly didn't know what to do, but now with your help i feel like i can pass the test next week!!!! thanks again for everything!!!!

OpenStudy (jhannybean):

yaaaay!!